OpenStudy (anonymous):

Consider the function f(x)=xe^−4x, 0≤x≤2. This function has an absolute maximum value equal to ? which is attained at x= ?

OpenStudy (anonymous):

the absolute min is at 0 and 0 respectively.

OpenStudy (anonymous):

I got x=1/4 see attached

OpenStudy (anonymous):

f'=e^(-4x)-4x*e^(-4x)=0 1-4x=0 x=1/4

OpenStudy (anonymous):

if x=1/4 is max, put the value of x in original equation & find f(1/4)

OpenStudy (anonymous):

worked, thanks :]

OpenStudy (anonymous):

welcome:D